# Rigidity aspects of Penrose’s singularity theorem

### v2

@@ 2 Preliminaries
Let $\Sigma\subset V$ be a MOTS, and let $\{\Sigma_{t}\}_{|t|<\varepsilon}$ be a variation of $\Sigma$ determined by $\Sigma_{t}\colon x\mapsto\exp_{x}\big{(}t\phi(x)\nu(x)\big{)}$ for some $\phi\in C^{\infty}(\Sigma)$. Let $\theta^{+}(t)$ denote the outward null expansion of $\Sigma_{t}$. From [2], we have that
@@
Nevertheless, it is shown in [2, 3] that, as a consequence of the Krein-Rutman theorem, $L$ admits a principal eigenvalue $\lambda_{1}$, which is real and simple. Furthermore there exists an associated principal eigenfunction $\phi$ which can be chosen to be strictly positive. Moreover, $\lambda_{1}\geq 0$ (respectively, $\lambda_{1}>0$) if and only if there exists a positive $\psi\in C^{\infty}(\Sigma)$ such that $L(\psi)\geq 0$ (respectively, $L(\psi)>0$). In analogy with the minimal surface case, we say a MOTS $\Sigma$ is stable if $\lambda_{1}\geq 0$.
The following lemma has been used in various rigidity results for MOTS. See [15, 17, 10, 20, 28].
@@
The following proposition is proved in the appendix of [11] and relies on the null maximum principle [14].
Proposition 2.
Let $(M,g)$ be a globally hyperbolic spacetime satisfying the null energy condition with a spacelike Cauchy hypersurface $V$. Let $\Sigma$ be a connected, closed hypersurface in $V$ which separates $V$. Suppose $\Sigma$ is a MOTS ($\theta^{+}=0$) with respect to the future directed null normal $\ell^{+}$ which points toward a component of $V\setminus\Sigma$ whose closure is noncompact. Then either $M$ is future null geodesically incomplete or else the future inextendible null normal geodesics in the direction of $\ell^{+}$ form an achronal totally geodesic ($\chi^{+}=0$) null hypersurface.
The next theorem can be thought of as a ‘one-sided version’ of Penrose’s singularity theorem, and it’s proved in a similar way [16].
@@ 3 Local rigidity
    there is a neighborhood $U$ of $\Sigma$ in $V$ diffeomorphic to $(-\varepsilon,\varepsilon)\times\Sigma$ such that for each $t\in(-\varepsilon,\varepsilon)$, $\Sigma_{t}=\{t\}\times\Sigma$ is a MOTS and the outward null normal geodesics to $\Sigma_{t}$ (past and future complete) form a closed achronal totally geodesic ($\chi^{+}_{t}=0$) null hypersurface $N_{t}$.
@@
Suppose $M$ is null geodesically complete. If there is a point on $\Sigma$ where $\theta^{+}<0$, then, by a result of Andersson and Metzger [4, Lem. 5.2], $\Sigma$ can be deformed via null mean curvature flow to an outer trapped surface $\Sigma^{\prime}$ which contradicts null geodesic completeness by Theorem 3. Thus $\Sigma$ is a MOTS.
@@
Therefore $\Sigma_{t}$ is inner trapped if we reverse the time orientation. Since the component in the inward direction also has noncompact closure, this contradicts past null completeness. Thus $\theta^{+}(t)=0$ for all $t\in(-\varepsilon,\varepsilon)$, i.e., $\Sigma_{t}$ a MOTS for all $t$. Proposition 2 along with its time-dual produce two totally geodesic achronal null hypersurfaces with boundary given by $\Sigma_{t}$; one is formed to the future, call it $F_{t}$, and the other formed to the past, call it $P_{t}$. We glue $F_{t}$ and $P_{t}$ together along $\Sigma_{t}$ forming a closed, totally geodesic null hypersurface $N_{t}$. Achronality of $N_{t}$ follows since $\Sigma_{t}$ separates $V$. Indeed, if $\gamma$ is a timelike curve starting on $P_{t}$ and ending on $F_{t}$, then the integral curves of a timelike vector field project $\gamma$ down to a curve in $V$ which must intersect $\Sigma_{t}$ at some point $x$. By concatenating this integral curve from $x$ with a portion of $\gamma$, we obtain a contradiction either with the achronality of $F_{t}$ or the achronality of $P_{t}$. Lastly, achronality of $N_{t}$ implies that $N_{t}=\partial I^{-}(\eta_{t})$ where $\eta_{t}$ is a null line in $M$ passing through $\Sigma_{t}$. ∎
@@
    $\theta^{\pm}=0$ and there are neighborhoods $U^{\pm}\approx(-\varepsilon^{\pm},\varepsilon^{\pm})\times\Sigma$ of $\Sigma$ in $V$ such that $U^{+}$ and $U^{-}$ are foliated by MOTS ($\theta^{+}=0$) and MITS ($\theta^{-}=0$), respectively. Each MOTS, $\Sigma^{+}_{t}$, generates a closed achronal totally geodesic null hypersurface, $N^{+}_{t}$, via the generators of the outward null normal $\ell^{+}_{t}$. Likewise, each MITS, $\Sigma^{-}_{s}$, generates an analogous $N^{-}_{s}$. Moreover, the intersection of the two families, $N^{+}_{t}$ and $N^{-}_{s}$, foliate a spacetime neighborhood of $\Sigma$ by totally geodesic codimension 2 spacelike submanifolds, each diffeomorphic to $\Sigma$.
@@
Suppose $M$ is null geodesically complete. We are assuming $\Sigma$ is weakly outer and inner trapped with respect to one of the components of $V\setminus\Sigma$. As in the proof of Theorem 4, $\Sigma$ is both a MOTS and a MITS. There are neighborhoods $U^{+}$ and $U^{-}$ of $\Sigma$ in $V$ foliated by MOTS and MITS, respectively. Call these foliations $U^{+}=\{\Sigma_{t}^{+}\}$ and $U^{-}=\{\Sigma_{s}^{-}\}$ for $t\in(-\varepsilon^{+},\varepsilon^{+})$ and $s\in(-\varepsilon^{-},\varepsilon^{-})$, and note that they coincide at $\Sigma=\Sigma_{0}^{+}=\Sigma_{0}^{-}$. Each $\Sigma_{t}^{+}$ generates a totally geodesic outward null hypersurface $N^{+}_{t}$ with vanishing second fundamental form $\chi^{+}_{t}=0$. Likewise with the inner direction: $\chi_{s}^{-}=0$ on $N_{s}^{-}$.
Let $N$ be the union of the intersections $\{N^{+}_{t}\cap N^{-}_{s}\mid t\in(-\varepsilon^{+},\varepsilon^{+}),\>s\in(-\varepsilon^{-},\varepsilon^{-})\}$. Then $N$ is nonempty since $\Sigma\subset N$. Since the null generators of these null hypersurfaces are distinct, it follows that $N^{+}_{t}$ and $N^{-}_{s}$ must intersect transversely (if they do at all). It follows that $N$ is an open set in $M$ containing $\Sigma$. Clearly, since $N^{+}_{0}\cap N^{-}_{0}=\Sigma$, $N^{+}_{t}\cap N^{-}_{s}$ will intersect in a bounded neighborhood of $\Sigma$ for all $t$ and $s$ sufficiently small in magnitude. The intersections within this bounded neighborhood are codimension 2 compact submanifolds which are totally geodesic since $\chi_{t}^{+}=\chi_{s}^{-}=0$. The null generators on $N^{+}_{t}$ establish a diffeomorphism from each intersection $N^{+}_{t}\cap N^{-}_{s}$ to $\Sigma^{+}_{t}$ and hence they are diffeomorphic to $\Sigma$. ∎
@@ 4 Compact-with-boundary rigidity
    $S$ is diffeomorphic to $\Sigma$ and $W$ is diffeomorphic to $[0,\delta]\times\Sigma$. Moreover, for each $t\in[0,\delta]$, $\Sigma_{t}=\{t\}\times\Sigma$ is a MOTS and the outward null normal geodesics to $\Sigma_{t}$ form a closed achronal totally geodesic ($\chi^{+}_{t}=0$) null hypersurface $N_{t}$.
@@
This follows from a continuity argument as in the proof of Theorem 1.2 in [10]. Assume $M$ is future null geodesically complete. Briefly, one shows that the leaves of the local foliation $\{\Sigma_{t}\}_{t\in[0,\varepsilon)}$ from Theorem 4 have a smooth embedded limit $\Sigma_{\varepsilon}$. (This uses the fact that, since $\chi_{t}=0$, the second fundamental forms of the $\Sigma_{t}$’s within $W$ are uniformly bounded in terms of the second fundamental form of $W$ in $M$.) Then the strong maximum principle ([6, Prop. 3.1], [4, Prop. 2.4]) shows that $\Sigma_{\varepsilon}=S$ if $\Sigma_{\varepsilon}\cap S\neq\emptyset$. If $\Sigma_{\varepsilon}\cap S=\emptyset$, then we can continue the foliation. Connectedness of $W$ implies $W\approx[0,\delta]\times\Sigma$ with $S\approx\Sigma_{\delta}$. ∎
Remark. Suppose in Theorem 6, one assumes that $V$ is geodesically complete, with bounded geometry, in the sense described in [27]. Suppose further that $K$ is uniformly bounded on all of $V$. Now let $W$ be the closure of one of the two components separating $V$. (Note that $W$ in this case is noncompact.) One can again argue that the leaves of the local foliation $\{\Sigma_{t}\}_{t\in[0,\varepsilon)}$ from Theorem 4 have a smooth embedded compact limit $\Sigma_{\varepsilon}$ provided the union of the leaves has a limit point in $W\setminus\Sigma$. Then, again by a continuation argument, if $M$ is future null geodesically complete, the conclusions (b) of Theorem 6 hold, with $[0,\delta]$ replaced by $[0,\delta)$, $\delta\in(0,\infty]$. We thank Michael Eichmair for clarifications regarding this.
@@
Remark. The second fundamental form $K$ on $V$ is said to be $(n-1)$-convex, if, at every point, the sum of the smallest $n-1$ eigenvalues of $K$ with respect to the induced metric $g|_{V}$ is nonnegative. It can be seen that this implies $\mathrm{tr}\,_{\Sigma}K\geq 0$ for every hypersurface $\Sigma\subset V$. If we assume in Theorem 7 that $K$ is $(n-1)$-convex, then the two foliations of MOTS $\Sigma^{+}_{t}$ and MITS $\Sigma^{-}_{s}$ actually agree with each other. To see this, fix $\Sigma^{-}_{s}$. We have $\theta^{-}_{s}=0$; hence $H=\text{tr}K\geq 0$ on $\Sigma^{-}_{s}$ by $(n-1)$-convexity. Therefore $\theta^{+}_{s}\geq 0$; hence $\theta^{+}_{s}=0$. By compactness, there is a $t$ such that $\Sigma^{+}_{t}$ and $\Sigma^{-}_{s}$ are tangent and $\Sigma^{-}_{s}$ lies to the inside of $\Sigma^{+}_{t}$. Since $\theta^{+}_{t}=\theta^{-}_{s}=0$, it follows that $\Sigma^{-}_{s}=\Sigma^{+}_{t}$ by the maximum principle for $\theta^{+}$ [6, Prop. 3.1].
@@ 6 Topological censorship rigidity
Topological censorship theorems show that the topology of the region exterior to a black hole is restricted in various respects, see e.g., [12, 13, 21, 26]. An important precursor to the principle of topological censorship (in conjunction with the weak cosmic censorship conjecture) is the Gannon-Lee singularity theorem [22, 25], which establishes conditions under which nontrivial topology leads to null geodesic incompleteness; see also [8, 9, 30, 19]. In this section we obtain some rigidity results in this context.
@@
    there is a neighborhood $U$ of $\Sigma$ in $V$ such that $U\approx(-\varepsilon,\varepsilon)\times\Sigma$ and for each $t\in(-\varepsilon,\varepsilon)$, $\Sigma_{t}=\{t\}\times\Sigma$ is a MITS and the inward null normal geodesics to $\Sigma_{t}$ form a closed totally geodesic ($\chi^{-}_{t}=0$) immersed null hypersurface.
@@
Suppose $M$ is future null geodesically complete. Set $E_{1}=\overline{V^{-}}$ and $E_{2}=\overline{V^{+}}$. Let $p\colon\widetilde{E}_{1}\to E_{1}$ denote the universal covering of $E_{1}$. By assumption $\widetilde{E}_{1}$ contains more than one copy of $\widetilde{\Sigma}:=p^{-1}(\Sigma)$. Since $\Sigma$ is simply connected, $\widetilde{\Sigma}$ is the disjoint union of $k$ diffeomorphic copies of $\Sigma$, where $k$ is the cardinality of $\pi_{1}(E_{1})$. Attach $k$ copies of $E_{2}$ to each copy of $\Sigma$ in $\widetilde{\Sigma}$ (as they’re attached in the base space). Call the resulting space $\widetilde{V}$, and, abusing notation, let $p\colon\widetilde{V}\to V$ denote the resulting covering. Using covering space theory (as outlined in the beginning of section 3.2 in [19]) or the Bernal-Sanchez splitting result (as outlined in [18]), there is a corresponding spacetime covering $P\colon\widetilde{M}\to M$ such that $\widetilde{V}$ is a Cauchy surface for $\widetilde{M}$ and $P|_{\widetilde{V}}=p$. Let $\hat{\Sigma}$ denote one of the components of $\widetilde{\Sigma}$. Since $k\geq 2$ and $E_{2}$ is noncompact, it follows that $\hat{\Sigma}$ separates $\widetilde{V}$ into two components whose closures are both noncompact. Therefore we can apply Theorem 4 (with outward and inward directions reversed) to $\hat{\Sigma}$ in the covering spacetime and then project down to the physical spacetime to obtain the result.∎
@@
Now assume $\overline{V^{-}}$ is not simply connected. The work of Hempel [24] in conjunction with the positive resolution of the geometrization conjecture shows that $\pi_{1}(\overline{V^{-}})$ is residually finite, i.e., for every non-identity element in the group, there is a normal subgroup with finite index that does not contain that element. (See [5] for a discussion.) Theorem 11 implies that $\overline{V^{-}}$ is double covered by $\Sigma\times[0,1]$. Again, attach a three-ball to $\overline{V^{-}}$ along $\Sigma$. The resulting space is double covered by $\mathbb{S}^{3}$ and hence it’s topologically $\mathbb{RP}^{3}$ by the positive resolution of the elliptization conjecture. The result follows. ∎

### v3

@@ 1 Introduction
Our results can be paraphrased in the following way “in the presence of a weakly trapped surface, either the spacetime is null geodesically complete (future and past) or the spacetime is very special in a suitably rigid sense.” Penrose’s original singularity theorem concluded only future null incompleteness, and our results can be interpreted in this spirit; see specifically Proposition 3 in section 2. Moreover, sections 4 and 6 are consistent with the setting of gravitational collapse developing from a nonsingular past. In this setting, our results imply that either the spacetime is future null geodesically incomplete or “the spacetime is very special in a suitably rigid sense.” A similar type of interpretation holds in the cosmological setting (section 5) where one naturally assumes future null completeness. In this setting, our results imply that either the spacetime is past null geodesically incomplete or “the spacetime is very special in a suitably rigid sense.”
@@ 2 Preliminaries
Let $\Sigma\subset V$ be a MOTS, and let $\{\Sigma_{t}\}_{|t|<\varepsilon}$ be a variation of $\Sigma$ determined by $\Sigma_{t}\colon x\mapsto\exp_{x}\big{(}t\phi(x)\nu(x)\big{)}$ for some $\phi\in C^{\infty}(\Sigma)$. Let $\theta^{+}(t)$ denote the outward null expansion of $\Sigma_{t}$. From e.g. [3, Eq. 1] (see also [26, Prop. 7.32]), we have that
@@
Nevertheless, it is shown in [3, 4] (see also [26, Thm. A.10]) that, as a consequence of the Krein-Rutman theorem, $L$ admits a principal eigenvalue $\lambda_{1}$, which is real and simple. Furthermore there exists an associated principal eigenfunction $\phi$ which can be chosen to be strictly positive. Moreover, $\lambda_{1}\geq 0$ (respectively, $\lambda_{1}>0$) if and only if there exists a positive $\psi\in C^{\infty}(\Sigma)$ such that $L(\psi)\geq 0$ (respectively, $L(\psi)>0$). In analogy with the minimal surface case, we say a MOTS $\Sigma$ is stable if $\lambda_{1}\geq 0$.
The following lemma has been used in various rigidity results for MOTS; see [16, 17, 11, 20, 29]. As it does not appear as a stand-alone result in the literature, we provide a proof in the appendix for the convenience of the reader.
@@
The next theorem can be thought of as a ‘one-sided version’ of Penrose’s singularity theorem, and it’s proved in a similar way, see e.g., [2, Thm. 7.1].
@@
The following proposition is a rigid version of the one-sided Penrose singularity theorem.
Proposition 3.
Let $(M,g)$ be a globally hyperbolic spacetime satisfying the null energy condition with a spacelike Cauchy hypersurface $V$. Let $\Sigma$ be a connected, closed hypersurface in $V$ which separates $V$. Suppose $\Sigma$ is a weakly trapped surface (i.e., $\theta^{+}\leq 0$) with respect to the future directed null normal $\ell^{+}$ which points toward a component of $V\setminus\Sigma$ whose closure is noncompact. Then either $M$ is future null geodesically incomplete or else $\Sigma$ is a MOTS (in fact $\chi^{+}=0$) and the future inextendible null normal geodesics in the direction of $\ell^{+}$ form an achronal totally geodesic [^3] By a totally geodesic null hypersurface, we mean that it’s null second fundamental form vanishes everywhere and has the usual interpretation that geodesics tangent to it remain within it. null hypersurface.
Proposition 3 is proved in the appendix of [12] in the case that $\Sigma$ is a MOTS ($\theta^{+}=0)$, but the proof trivially carries over to the case when $\Sigma$ is weakly trapped.
@@ 3 Local rigidity
    there is a neighborhood $U$ of $\Sigma$ in $V$ diffeomorphic to $(-\varepsilon,\varepsilon)\times\Sigma$ such that for each $t\in(-\varepsilon,\varepsilon)$, $\Sigma_{t}=\{t\}\times\Sigma$ is a MOTS (in fact $\chi^{+}_{t}=0)$ and the outward null normal geodesics to $\Sigma_{t}$ (past and future complete) form a closed achronal totally geodesic null hypersurface $N_{t}$.
@@
Suppose $M$ is null geodesically complete. If there is a point on $\Sigma$ where $\theta^{+}<0$, then, by a result of Andersson and Metzger [5, Lem. 5.2], $\Sigma$ can be deformed via null mean curvature flow within $V$ to an outer trapped surface $\Sigma^{\prime}$ which contradicts null geodesic completeness by Theorem 2. Thus $\Sigma$ is a MOTS.
@@
Therefore $\Sigma_{t}$ is inner trapped if we reverse the time orientation. Since the component in the inward direction also has noncompact closure, this contradicts past null completeness. Thus $\theta^{+}(t)=0$ for all $t\in(-\varepsilon,\varepsilon)$, i.e., $\Sigma_{t}$ a MOTS for all $t$. Proposition 3 along with its time-dual produce two totally geodesic achronal null hypersurfaces with boundary given by $\Sigma_{t}$; one is formed to the future, call it $F_{t}$, and the other formed to the past, call it $P_{t}$. We glue $F_{t}$ and $P_{t}$ together along $\Sigma_{t}$ forming a closed, totally geodesic null hypersurface $N_{t}$. Achronality of $N_{t}$ follows since $\Sigma_{t}$ separates $V$. Indeed, if $\gamma$ is a timelike curve starting on $P_{t}$ and ending on $F_{t}$, then the integral curves of a timelike vector field project $\gamma$ down to a curve in $V$ which must intersect $\Sigma_{t}$ at some point $x$. By concatenating this integral curve from $x$ with a portion of $\gamma$, we obtain a contradiction either with the achronality of $F_{t}$ or the achronality of $P_{t}$. ∎
@@
    $\theta^{\pm}=0$ and there are neighborhoods $U^{\pm}\approx(-\varepsilon^{\pm},\varepsilon^{\pm})\times\Sigma$ of $\Sigma$ in $V$ such that $U^{+}$ and $U^{-}$ are foliated by MOTS (in fact $\chi^{+}=0$) and MITS (in fact $\chi^{-}=0$), respectively. Each MOTS, $\Sigma^{+}_{t}$, generates a closed achronal totally geodesic null hypersurface, $N^{+}_{t}$, via the generators of the outward null normal $\ell^{+}_{t}$. Likewise, each MITS, $\Sigma^{-}_{s}$, generates an analogous $N^{-}_{s}$. Moreover, the intersection of the two families, $N^{+}_{t}$ and $N^{-}_{s}$, foliate a spacetime neighborhood of $\Sigma$ by totally geodesic codimension 2 spacelike submanifolds, each diffeomorphic to $\Sigma$.
@@
Suppose $M$ is null geodesically complete. We are assuming $\Sigma$ is weakly outer and inner trapped with respect to one of the components of $V\setminus\Sigma$. As in the proof of Theorem 4, $\Sigma$ is both a MOTS and a MITS. There are neighborhoods $U^{+}$ and $U^{-}$ of $\Sigma$ in $V$ foliated by MOTS and MITS, respectively. Call these foliations $U^{+}=\{\Sigma_{t}^{+}\}$ and $U^{-}=\{\Sigma_{s}^{-}\}$ for $t\in(-\varepsilon^{+},\varepsilon^{+})$ and $s\in(-\varepsilon^{-},\varepsilon^{-})$, and note that they coincide at $\Sigma=\Sigma_{0}^{+}=\Sigma_{0}^{-}$. Each $\Sigma_{t}^{+}$ generates a totally geodesic outward null hypersurface $N^{+}_{t}$, i.e., it has vanishing null second fundamental. Likewise with the inner direction on $N_{s}^{-}$.
Let $N$ be the union of the intersections $\{N^{+}_{t}\cap N^{-}_{s}\mid t\in(-\varepsilon^{+},\varepsilon^{+}),\>s\in(-\varepsilon^{-},\varepsilon^{-})\}$. Then $N$ is nonempty since $\Sigma\subset N$. Since the null generators of these null hypersurfaces are distinct, it follows that $N^{+}_{t}$ and $N^{-}_{s}$ must intersect transversely (if they do at all). It follows that $N$ is an open set in $M$ containing $\Sigma$. Clearly, since $N^{+}_{0}\cap N^{-}_{0}=\Sigma$, $N^{+}_{t}\cap N^{-}_{s}$ will intersect in a spacetime neighborhood of $\Sigma$ for all $t$ and $s$ sufficiently small in magnitude. The intersections within this neighborhood are codimension 2 compact submanifolds which are totally geodesic since it’s the intersection of two transverse totally geodesic null hypersurfaces. The null geodesic generators of $N^{+}_{t}$ establish a diffeomorphism from each intersection $N^{+}_{t}\cap N^{-}_{s}$ to $\Sigma^{+}_{t}$ and hence they are diffeomorphic to $\Sigma$. Indeed, by an achronality argument, the null geodesic generators of $N^{+}_{t}$ intersect $N^{-}_{s}$ only once. Therefore the exponential map sets up a well-defined embedding of $\Sigma^{+}_{t}$ into $N^{+}_{t}\cap N^{-}_{s}$ producing the diffeomorphism. ∎
@@ 4 Compact-with-boundary rigidity
    $S$ is diffeomorphic to $\Sigma$ and $W$ is diffeomorphic to $[0,\delta]\times\Sigma$. Moreover, for each $t\in[0,\delta]$, $\Sigma_{t}=\{t\}\times\Sigma$ is a MOTS (in fact $\chi^{+}_{t}=0$) and the outward null normal geodesics to $\Sigma_{t}$ form a closed achronal totally geodesic null hypersurface $N_{t}$.
@@
This follows from a continuity argument as in the proof of Theorem 1.2 in [11]. Assume $M$ is future null geodesically complete. Briefly, one shows that the leaves of the local foliation $\{\Sigma_{t}\}_{t\in[0,\varepsilon)}$ from Theorem 4 have a smooth embedded limit $\Sigma_{\varepsilon}$. (This uses the fact that, since $\chi_{t}^{+}=0$, the second fundamental forms of the $\Sigma_{t}$’s within $W$ are uniformly bounded in terms of the second fundamental form of $W$ in $M$.) Then the strong maximum principle ([7, Prop. 3.1], [5, Prop. 2.4]) shows that $\Sigma_{\varepsilon}=S$ if $\Sigma_{\varepsilon}\cap S\neq\emptyset$. If $\Sigma_{\varepsilon}\cap S=\emptyset$, then we can continue the foliation. Connectedness of $W$ implies $W\approx[0,\delta]\times\Sigma$ with $S\approx\Sigma_{\delta}$. ∎
Remark. Suppose in Theorem 6, one assumes that $V$ is geodesically complete, with bounded geometry, in the sense described in [28, Sec. 1] (called ‘homogeneous regularity’ there). Suppose further that $K$ is uniformly bounded on all of $V$. Now let $W$ be the closure of one of the two components separating $V$. (Note that $W$ in this case is noncompact.) One can again argue that the leaves of the local foliation $\{\Sigma_{t}\}_{t\in[0,\varepsilon)}$ from Theorem 4 have a smooth embedded compact limit $\Sigma_{\varepsilon}$ provided the union of the leaves has a limit point in $W\setminus\Sigma$. Then, again by a continuation argument, if $M$ is future null geodesically complete, the conclusions (b) of Theorem 6 hold, with $[0,\delta]$ replaced by $[0,\delta)$, $\delta\in(0,\infty]$. We thank Michael Eichmair for clarifications regarding this.
In the following, we apply the previous theorem to a situation where $\Sigma$ is both weakly outer and inner trapped. A priori, the null expansion assumptions on $S$ are compatible with an asymptotically flat end.
@@
Remark. The second fundamental form $K$ on $V$ is said to be $(n-1)$-convex, if, at every point, the sum of the smallest $n-1$ eigenvalues of $K$ with respect to the induced metric $g|_{V}$ is nonnegative. It can be seen that this implies $\mathrm{tr}\,_{\Sigma}K\geq 0$ for every hypersurface $\Sigma\subset V$. If we assume in Theorem 7 that $K$ is $(n-1)$-convex, then the two foliations of MOTS $\Sigma^{+}_{t}$ and MITS $\Sigma^{-}_{s}$ actually agree with each other. To see this, fix $\Sigma^{-}_{s}$. We have $\theta^{-}_{s}=0$; hence $H=\text{tr}K\geq 0$ on $\Sigma^{-}_{s}$ by $(n-1)$-convexity. Therefore $\theta^{+}_{s}\geq 0$; hence $\theta^{+}_{s}=0$. By compactness, there is a $t$ such that $\Sigma^{+}_{t}$ and $\Sigma^{-}_{s}$ are tangent and $\Sigma^{-}_{s}$ lies to the inside of $\Sigma^{+}_{t}$. Since $\theta^{+}_{t}=\theta^{-}_{s}=0$, it follows that $\Sigma^{-}_{s}=\Sigma^{+}_{t}$ by the maximum principle for $\theta^{+}$ [7, Prop. 3.1].
@@ 6 Topological censorship rigidity
Topological censorship theorems show that the topology of the region exterior to a black hole is restricted in various respects, see e.g., [13, 14, 21, 27]. An important precursor to the principle of topological censorship (in conjunction with the weak cosmic censorship conjecture) is the Gannon-Lee singularity theorem ([22, Prop. 1.2], [25]), which establishes conditions under which nontrivial topology leads to null geodesic incompleteness; see also [9, 10, 31, 19]. In this section we obtain some rigidity results in this context.
@@
    there is a neighborhood $U$ of $\Sigma$ in $V$ such that $U\approx(-\varepsilon,\varepsilon)\times\Sigma$ and for each $t\in(-\varepsilon,\varepsilon)$, $\Sigma_{t}=\{t\}\times\Sigma$ is a MITS (in fact $\chi^{-}_{t}=0$) and the inward null normal geodesics to $\Sigma_{t}$ form a closed totally geodesic immersed null hypersurface.
@@
Suppose $M$ is future null geodesically complete. Set $E_{1}=\overline{V^{-}}$ and $E_{2}=\overline{V^{+}}$. Let $p\colon\widetilde{E}_{1}\to E_{1}$ denote the universal covering of $E_{1}$. By assumption $\widetilde{E}_{1}$ contains more than one copy of $\widetilde{\Sigma}:=p^{-1}(\Sigma)$. Since $\Sigma$ is simply connected, $\widetilde{\Sigma}$ is the disjoint union of $k$ diffeomorphic copies of $\Sigma$, where $k$ is the cardinality of $\pi_{1}(E_{1})$. Attach $k$ copies of $E_{2}$ to each copy of $\Sigma$ in $\widetilde{\Sigma}$ (as they’re attached in the base space). Call the resulting space $\widetilde{V}$, and, abusing notation, let $p\colon\widetilde{V}\to V$ denote the resulting covering. Using covering space theory (as outlined in the beginning of section 3.2 in [19]) or the Bernal-Sanchez splitting result (as outlined in [18, Lemma 4]), there is a corresponding spacetime covering $P\colon\widetilde{M}\to M$ such that $\widetilde{V}$ is a Cauchy surface for $\widetilde{M}$ and $P|_{\widetilde{V}}=p$. Let $\hat{\Sigma}$ denote one of the components of $\widetilde{\Sigma}$. Since $k\geq 2$ and $E_{2}$ is noncompact, it follows that $\hat{\Sigma}$ separates $\widetilde{V}$ into two components whose closures are both noncompact. Therefore we can apply Theorem 4 (with outward and inward directions reversed) to $\hat{\Sigma}$ in the covering spacetime and then project down to the physical spacetime to obtain the result.∎
@@
Now assume $\overline{V^{-}}$ is not simply connected. The work of Hempel [24, Thm. 1.1] in conjunction with the positive resolution of the geometrization conjecture shows that $\pi_{1}(\overline{V^{-}})$ is residually finite, i.e., for every non-identity element in the group, there is a normal subgroup with finite index that does not contain that element. (See the introduction of [6] for a discussion.) Theorem 11 implies that $\overline{V^{-}}$ is double covered by $\Sigma\times[0,1]$. Again, attach a three-ball to $\overline{V^{-}}$ along $\Sigma$; in the cover this corresponds to capping off both ends of $\Sigma\times[0,1]$ by three-balls to obtain a three-sphere. Hence the resulting space is double covered by $\mathbb{S}^{3}$ and hence it’s topologically $\mathbb{RP}^{3}$ by the positive resolution of the elliptization conjecture. The result follows. ∎
@@
Appendix A Proof of Lemma 1
Proof of Lemma 1. For small $f\in C^{\infty}(\Sigma)$, let $\Theta^{+}(f)$ denote the outward null expansion of the hypersurface $\Sigma_{f}\colon x\to\exp_{x}\big{(}f(x)\nu\big{)}$ with respect to the (suitably normalized) future directed outward null normal field to $\Sigma_{u}$. $\Theta^{+}$ has linearization, $\Theta^{+\prime}(0)=L$, where $L$ is the MOTS stability operator introduced in section 2. We introduce the operator
$\Phi\colon C^{\infty}(\Sigma)\times\mathbb{R}\to C^{\infty}(\Sigma)\times\mathbb{R},\quad\Phi(f,k)\,=\,\left(\Theta^{+}(f)-k,\int_{\Sigma}f\right).$
Since $\lambda_{1}=0$ is a simple eigenvalue, the kernel of $\Theta^{+\prime}(0)=L$ consists only of constant multiples of the positive eigenfunction $\phi$. Moreover, $\lambda_{1}=0$ is a simple eigenvalue for the adjoint $L^{*}$ of $L$ [4, Lem. 4.1(ii)] (with respect to the $L^{2}$ inner product on $\Sigma$); let $\phi^{*}$ denote a corresponding positive eigenfunction. Then the equation $Lf=v$ is solvable if and only if $\int_{\Sigma}v\phi^{*}=0$ (by the Fredholm alternative for elliptic operators). The linearization of $\Phi$ is
$(f,k)\mapsto\left(Lf-k,\int_{\Sigma}f\right).$
We show $\Phi$ has invertible linearization about $(0,0)$. Injectivity: Suppose $Lf-k=0$ and $\int_{\Sigma}f=0$; the former implies $\int_{\Sigma}k\phi^{*}=0$ which implies $k=0$; hence $f$ is a multiple of $\phi$ and so the latter implies $f=0$. Surjectivity: Fix a function $h\in C^{\infty}(\Sigma)$ and $a\in\mathbb{R}$. We want to find a function $f$ and number $k$ such that $Lf-k=h$ and $\int_{\Sigma}f=a$. $Lf=h+k$ implies $\int_{\Sigma}\phi^{*}(h+k)=0$; thus we must have $k=-(\int_{\Sigma}\phi^{*}h)/(\int_{\Sigma}\phi^{*})$. There exists an $f_{0}$ such that $Lf_{0}=h+k$. Consider $f=f_{0}+c\phi$. Then $Lf=h+k$ also. Moreover, $\int_{\Sigma}f=\int_{\Sigma}f_{0}+c\int_{\Sigma}\phi$. Therefore choose $c$ such that $\int_{\Sigma}f=a$.
Thus, by the inverse function theorem, for $t\in\mathbb{R}$ sufficiently small, there exists $f(t)\in C^{\infty}(\Sigma)$ and $k(t)\in\mathbb{R}$ such that
$\Theta^{+}\big{(}f(t)\big{)}\,=\,k(t)\quad\text{ and }\quad\int_{\Sigma}f(t)\,=\,t.$
By the chain rule, we have $\Theta^{+\prime}(0)\big{(}f^{\prime}(0)\big{)}=L\big{(}f^{\prime}(0)\big{)}=k^{\prime}(0)$. Therefore $k^{\prime}(0)$ is orthogonal to $\phi^{*}$. Hence $k^{\prime}(0)=0$, and so $f^{\prime}(0)$ is in the kernel of $L$. Thus $f^{\prime}(0)$ is a constant multiple of $\phi$, and since $\int_{\Sigma}f^{\prime}(0)=1$, it follows that $f^{\prime}(0)>0$. Moreover, by invertibility, we have $f(0)=0$.
It follows that for $t$ sufficiently small, the hypersurfaces $\Sigma_{t}:=\Sigma_{f(t)}$ form a smooth foliation of a neighborhood of $\Sigma$ in $V$ by hypersurfaces of constant outward null expansion. Thus, one can introduce coordinates $(t,x^{i})$ in a neighborhood $U$ of $\Sigma$ in $V$ such that, with respect to these coordinates, $U=(-\varepsilon,\varepsilon)\times\Sigma$, and for each $t\in(-\varepsilon,\varepsilon)$, the $t$-slice $\Sigma_{t}=\{t\}\times S$ has constant outward null expansion $\theta^{+}(t)$ with respect to $\ell(t)=u+\nu(t)$ where $\nu$ is the outward unit normal field to $\Sigma_{t}$ in $V$. Moreover, the coordinates $(t,x^{i})$ can be chosen so that $\partial_{t}=\phi\nu$, for some positive function $\phi$ on $U$. It follows that the metric on $U$ is given by $h|_{U}=\phi^{2}dt^{2}+\gamma_{t}$ where $\gamma_{t}$ is the induced metric on $\Sigma_{t}$. ∎